In a right triangle, find the length of the side not given.
b=2
c=sqrt29
a=?

Answer by Alan3354(30924)

In a right triangle, find the length of the side not given.
b=2
c=sqrt29
a=?
-------------
Side c could be the hypotenuse, or another shorter side.
----
If it's the hypotenuse:
a^2 = c^2 - b^2
a^2 = 29 - 4
a = 5
-----------
If it's not the hypotenuse:
a^2 = c^2 + b^2
a^2 = 29 + 4
a = sqrt(33)